if the total amount of fencing that the farmer has is 20 metres,write an equation using the lengthand breath of the garden
Answer by london maths tutor(243)

Let's length be "L" and breath be "b"
Perimetre = 2L + 2b = 2(L+b)
Therefore equation = 2(L+b) = 20
Simplify further : L + b = 10 (Answer)